Exercise 16-1

The CPA Partnership operated by Cook, Parks, and Argo is being liquidated. A balance sheet prepared at this stage in their liquidation process is presented below.
Cash $44,000 Liabilities $34,000
Other Assets 51,000 Parks, Loan 10,000
Cook, Capital 28,000
Parks, Capital 9,000
Argo, Capital 14,000
   Total $95,000    Total $95,000

The partners share profits and losses 30% (Cook), 50% (Parks), and 20% (Argo). The partners are all personally insolvent.

Collapse question part

(a)

The partners wish to distribute the $44,000 in cash. Record in journal entry form the distribution of the available cash. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit

(To record payment of liabilities)

(To record distribution to partners)

SHOW LIST OF ACCOUNTS
Answer #1

Solution: (Amounts in $)

Particulars Debit Credit

1. For Payment of Liabilities

Liabilities A/c Dr.

To Cash A/c

(Being liabilities paid fully)

34,000

34,000

2. For distribution of remaining cash to all partners

Cook Capital A/c Dr.

Parks Capital A/c Dr.

Argo Capital A/c Dr.

To Cash A/c

(Being remaining cash distributed to all partners in their capital ratios)

5,491

1,765

2,744

10,000

Working Note: Distribution of remaining cash to all partners in their capital ratios:

Total cash balance = $ 44,000

Less: Paid to liabilities = ($ 34,000)

Available balance = $ 10,000

Distribution:

Cook Capital A/c = $ 10,000 x 28/51 = $ 5,491

Parks Capital A/c = $ 10,000 x 9/51 = $ 1,765

Argo Capital A/c = $ 10,000 x 14/51 = $ 2,744